The radius of a quarter circle is 7 meters. What is the quarter circle's perimeter?


Use 3.14 for ​𝜋.


1
Expert's answer
2022-02-17T03:24:00-0500

Arc length of quarter circle "=\frac{\theta}{360^0}2\pi r=\frac{90^0}{360^0} \times2\times 3.14 \times 7=10.99m"

quarter circle's perimeter "=10.99+7+7=24.99m"